OpenStudy (anonymous):

What is the exact value of the expressions the square root of 180. - the square root 125. + the square root of 5.? Simplify if possible.

OpenStudy (anonymous):

@Jhannybean

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

\[\sqrt{180}-\sqrt{125}-\sqrt{5}\] \[=\sqrt{36*5}-\sqrt{25*5}-\sqrt{5}=?\]

OpenStudy (anonymous):

WHATS THE ANSWER

OpenStudy (anonymous):

HOW DO U ANSWER IT I WILL GIV A MEDAL

OpenStudy (anonymous):

\[=6\sqrt{5}-5\sqrt{5}-\sqrt{5}=6\sqrt{5}-6\sqrt{5}=?\]